OpenStudy (anonymous):

6-4(3)______5-10 A.< B.> C.= D.none of the above

OpenStudy (anonymous):

what is 4(3)?

OpenStudy (anonymous):

idk ? that is what i am have a hard time figuring it out

OpenStudy (anonymous):

6-12=? and 5-10=?

OpenStudy (anonymous):

4(3) means 4*3

OpenStudy (anonymous):

ok that makes sense

OpenStudy (anonymous):

so it would be B

OpenStudy (anonymous):

so solve for this 6-12=? and 5-10=? and you get -6______-5 so what sign would go in there

OpenStudy (anonymous):

although the number may be bigger if you look on a number line -6 is smaller then -5 so it wouldnt be B

OpenStudy (anonymous):

it would be A

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np